OpenStudy (anonymous):

To compose music on a computer, you can write notes as decimals. What decimals should you enter for a half not, a quarter note, an eighth note, and a sixteenth note?

OpenStudy (anonymous):

This is simple no? What is a half? 1/2 = .5 A quarter? Four quarters make a whole. So one is 1/4 = .25 The rest should be straight forward. ;)

OpenStudy (anonymous):

This is simple no? What is a half? 1/2 = .5 A quarter? Four quarters make a whole. So one is 1/4 = .25 The rest should be straight forward. ;)

OpenStudy (anonymous):

This is simple no? What is a half? 1/2 = .5 A quarter? Four quarters make a whole. So one is 1/4 = .25 The rest should be straight forward. ;)